# Long-Tailed Classification Based on Coarse-Grained Leading Forest and Multi-center Loss

### v2

@@ IV Experiment
TABLE I: Evaluation of CLT and GLT protocols on ImageNet-GLT. Among 48 comparisons against the competing models, the family of ?+Cognisance won for 43 times.
@@
Re-balance Baseline width 1pt $58.39|38.35$ $36|52.15$ $13.98|55.34$ $41.65|47.11$ width 1pt $50.7|32.5$ $27.8|43.99$ $10.18|47.7$ $34.32|39.95$
BBN width 1pt $62.66|43.45$ $44.36|55.44$ $14.66|57.57$ $47.22|50.96$ width 1pt $53.37|36.45$ $34.97|46.81$ $10.73|46.08$ $38.7|42.56$
@@
Logit-Adj width 1pt $53.3|49.05$ $43.49|44.12$ $31.86|35$ $45.67|44.72$ width 1pt $45.43|41.75$ $34.2|35.9$ $24.58|27.78$ $37.25|37.02$
GLTv1 width 1pt $62.51|42.32$ $38.69|57.23$ $17.41|\textbf{65}$ $45.03|52.43$ width 1pt $54.4|36.23$ $30.18|49.61$ $12.64|55.47$ $37.24|45.14$
@@
* BLSoftmax + Cognisance width 1pt $58.37|53.88$ $\textbf{44.98}|51.97$ $33.82|39.05$ $\textbf{48.66}|50.8$ width 1pt $49.92|46.89$ $\textbf{36.11}|44.31$ $26.14|30.03$ $\textbf{40.14}|43.2$
* Logit-Adj + Cognisance width 1pt $43.3|\textbf{75.13}$ $41.19|\textbf{59.08}$ $\textbf{45.43}|27.11$ $42.92|\textbf{60.7}$ width 1pt $36.01|\textbf{70.18}$ $32.6|\textbf{52.48}$ $\textbf{36.26}|21.76$ $34.51|\textbf{54.95}$
@@
RandAug width 1pt $64.14|42.23$ $40.1|58.27$ $14.96|59.51$ $45.94|52.04$ width 1pt $55.7|35.87$ $31.61|50.15$ $10.2|47.29$ $38.03|44.01$
@@
* RandAug + Cognisance width 1pt $\textbf{69.12}|\textbf{49.31}$ $\textbf{47.64}|\textbf{63.08}$ $\textbf{26.97}|67.01$ $\textbf{53.13}|\textbf{58.16}$ width 1pt $\textbf{60.85}|\textbf{42.3}$ $\textbf{38.46}|\textbf{55.21}$ $\textbf{19.9}|\textbf{60.11}$ $\textbf{44.63}|\textbf{50.78}$
Ensemble TADE width 1pt $57.3|\textbf{55.22}$ $46.85|50.29$ $\textbf{34.69}|37.93$ $49.21|50.41$ width 1pt $49.61|\textbf{48.19}$ $37.55|42.59$ $\textbf{27.52}|32.21$ $40.87|43.28$
RIDE width 1pt $63.18|51.44$ $47.67|52.55$ $29.91|47.38$ $51.21|51.33$ width 1pt $54.83|44.02$ $38.25|44.2$ $22.77|38.12$ $42.56|43.21$
@@
* RIDE + Cognisance width 1pt $\textbf{64.83}|54.6$ $\textbf{50.95}|\textbf{56.21}$ $33.28|\textbf{50.15}$ $\textbf{53.85}|\textbf{54.66}$ width 1pt $\textbf{57.1}|47.63$ $\textbf{42}|\textbf{48.42}$ $25.5|\textbf{41.23}$ $\textbf{45.57}|{47.03}$
TABLE II: Evaluation of CLT and GLT protocols on MSCOCO-GLT. Among 48 comparisons against the competing models, the family of ?+Cognisance won for 44 times.
@@
Re-balance Baseline width 1pt $81.27|71.08$ $74.13|76.06$ $50.17|85.61$ $71.88|76.15$ width 1pt $74.59|64.82$ $66.25|69.08$ $35.75|77.22$ $63.1|69.14$
BBN width 1pt $83.59|70.21$ $76.13|76.3$ $47.25|\textbf{90.98}$ $72.98|77.02$ width 1pt $75.36|61.97$ $68.29|68.61$ $31.75|\textbf{81.38}$ $63.41|68.73$
BLSoftmax width 1pt $80.77|71.87$ $75.67|69.94$ $45.25|90.08$ $71.31|74.84$ width 1pt $73.23|65.13$ $68.63|62.64$ $32.08|79.6$ $62.81|67.09$
Logit-Adj width 1pt $81.55|73.95$ $76|75.55$ $\textbf{60.83}|82.04$ $74.97|76.28$ width 1pt $73.64|66.98$ $68.71|67.61$ $\textbf{46.58}|70.69$ $66|68.01$
GLTv1 width 1pt $82.45|73.09$ $76.42|79.53$ $55.58|86.87$ $74.4|78.6$ width 1pt $76.41|67.07$ $67.33|70.54$ $39.75|81.02$ $65.07|71.39$
@@
* BLSoftmax + Cognisance width 1pt $83|74.17$ $78.37|76.24$ $53.92|86.51$ $75.07|77.58$ width 1pt $75.86|67.98$ $70.46|68.14$ $40.08|80.26$ $66.22|70.59$
* Logit-Adj + Cognisance width 1pt $82.68|\textbf{77.16}$ $\textbf{80.17}|75.45$ $59.17|83.54$ $\textbf{76.78}|77.77$ width 1pt $75.14|\textbf{70.56}$ $\textbf{71.54}|66.34$ $45.83|73.9$ $\textbf{67.59}|69.5$
Augment Mixup width 1pt $82.41|72.95$ $73.12|79.53$ $54.33|87.47$ $72.76|78.68$ width 1pt $74.64|66.43$ $65.33|71$ $36|78.11$ $62.79|70.74$
@@
* RandAug + Cognisance width 1pt $\textbf{86.55}|\textbf{78.31}$ $81.5|82.82$ $62.42|\textbf{90.29}$ $\textbf{79.47}|\textbf{82.65}$ width 1pt $\textbf{79.05}|69.93$ $\textbf{72.92}|74.51$ $44.33|\textbf{81.82}$ $69.33|74.28$
Ensemble TADE width 1pt $83.55|76.47$ $80.79|73.89$ $50.5|89.34$ $75.57|78.06$ width 1pt $77.18|68.62$ $71.58|65.63$ $33.5|82.32$ $65.83|70.22$
RIDE width 1pt $84.23|77.56$ $82.29|77.79$ $58.42|89.69$ $78.09|80.16$ width 1pt $77.45|71.2$ $75.17|68.67$ $41.25|83.33$ $69.02|72.66$
@@
* RIDE + Cognisance width 1pt $\textbf{86.27}|\textbf{79.64}$ $\textbf{83.13}|\textbf{79.89}$ $\textbf{63.5}|\textbf{90.41}$ $\textbf{80.26}|\textbf{81.98}$ width 1pt $\textbf{80.09}|\textbf{72.77}$ $\textbf{75.17}|\textbf{71.45}$ $\textbf{47.83}|83.03$ $\textbf{71.38}|\textbf{74.34}$
IV-A Evaluation Protocols
  • Class wise Long Tail (CLT) Protocol: The samples in the training set follow a long-tailed distribution, which means that they are normally sampled from the LT dataset, while the samples in the test set are balanced by class. Note that the issue of attribute distribution is not considered in CLT, as the training and testing sets of CLT have the same attribute distribution and different class distributions, so the effectiveness of category long tail classification can be evaluated.
IV-B Datasets and Metrics
@@ IV-C Comparisons with LT Line-up
TABLE III: Evaluation of ImageNet-GLT and MSCOCO-GLT on LT test set. Among 18 comparisons against the competing models, the family of ?+Cognisance won for 16 times.
Benchmarks width 1pt ImageNet-GLT width 1pt MSCOCO-GLT
Overall Evaluation width 1pt $Acc$ $Prec$ $F1$ width 1pt $Acc$ $Prec$ $F1$
BBN width 1pt 58.6 48.9 53.32 width 1pt 84.84 78.04 81.3
BLSoftmax width 1pt 51.73 41.97 46.34 width 1pt 83.69 73.81 78.44
Logit-Adj width 1pt 50.94 40.58 45.17 width 1pt 85.28 73.6 79.01
GLTv1 width 1pt 58.28 50.43 54.07 width 1pt 86.67 81.88 84.21
@@
Augment Mixup width 1pt 55.25 46.76 50.66 width 1pt 86.95 82 84.4
RandAug width 1pt 59.88 51.28 55.25 width 1pt 87.64 81.02 84.2
@@
Ensemble TADE width 1pt 55.21 46.52 50.49 width 1pt 86.57 79.91 83.1
RIDE width 1pt 60.17 48.28 53.58 width 1pt 88.03 81.79 84.8
@@
* RIDE + Cognisance width 1pt 62.11 50.8 55.89 width 1pt 89.02 81.85 85.28
@@ V-A Why not directly apply the OLeaF?
In Cognisance, CLF is more appropriate compared to OLeaF for two very important characteristics: 1) Automatic tree-splitting mechanism. The number of parameters of CLF is relatively small, and the fine degree of tree-splitting can be controlled only by adjusting the parameter $d_{rd}$. The tree-splitting scheme in OLeaF is more rigorous and detailed but requires manually tuning parameters $d_{c}$ and number of trees. However, the feature learner needs to iterate in this framework and it may be possible to re-trigger clustering after each iteration. Besides, the target datasets for Cognisance are often large image data, so a fully automated scheme is compulsory; 2) Computing efficiency and representativeness. Each node of CLF is a Coarse-Grained node because the head attribute may contain a large number of similar samples. If left un-merged, the length of the path occupied by the head attribute may skyrocket. While each node in the same path is sampled with the same weight, a large number of extremely similar nodes actually lose their representativeness. At the same time, the sampling weights of bottom nodes are compromised. So it is necessary to carry out a small range of clustering to be controlled, and this clustering process is mainly controlled by the parameter $d_{rn}$, and all the member nodes within a coarse node will equally share the node’s sampling weight.

### v3

@@ IV Experiment
IV-A Evaluation Protocols
  • Class wise Long Tail (CLT) Protocol: The samples in the training set follow a long-tailed distribution, which means that they are normally sampled from the LT dataset, while the samples in the test set are balanced by class. Note that the issue of attribute distribution is not considered in CLT, as the training and testing sets of CLT have the same attribute distribution and different class distributions, so the effectiveness of category long-tailed classification can be evaluated.
IV-B Datasets and Metrics
TABLE I: Evaluation of CLT and GLT protocols on ImageNet-GLT. The family of ?+Cognisance won for 43 times out of 48 comparisons against the competing models.
@@ IV-B Datasets and Metrics
Re-balance Baseline   width 1pt $58.39|38.35$ $36.00|52.15$ $13.98|55.34$ $41.65|47.11$   width 1pt $50.70|32.50$ $27.80|43.99$ $10.18|47.70$ $34.32|39.95$
BBN   width 1pt $62.66|43.45$ $44.36|55.44$ $14.66|57.57$ $47.22|50.96$   width 1pt $53.37|36.45$ $34.97|46.81$ $10.73|46.08$ $38.70|42.56$
@@
Logit-Adj   width 1pt $53.30|49.05$ $43.49|44.12$ $31.86|35.00$ $45.67|44.72$   width 1pt $45.43|41.75$ $34.20|35.90$ $24.58|27.78$ $37.25|37.02$
GLTv1   width 1pt $62.51|42.32$ $38.69|57.23$ $17.41|\textbf{65.00}$ $45.03|52.43$   width 1pt $54.40|36.23$ $30.18|49.61$ $12.64|55.47$ $37.24|45.14$
@@
* BLSoftmax + Cognisance   width 1pt $58.37|53.88$ $\textbf{44.98}|51.97$ $33.82|39.05$ $\textbf{48.66}|50.80$   width 1pt $49.92|46.89$ $\textbf{36.11}|44.31$ $26.14|30.03$ $\textbf{40.14}|43.20$
* Logit-Adj + Cognisance   width 1pt $43.30|\textbf{75.13}$ $41.19|\textbf{59.08}$ $\textbf{45.43}|27.11$ $42.92|\textbf{60.70}$   width 1pt $36.01|\textbf{70.18}$ $32.60|\textbf{52.48}$ $\textbf{36.26}|21.76$ $34.51|\textbf{54.95}$
@@
RandAug   width 1pt $64.14|42.23$ $40.10|58.27$ $14.96|59.51$ $45.94|52.04$   width 1pt $55.70|35.87$ $31.61|50.15$ $10.20|47.29$ $38.03|44.01$
@@
* RandAug + Cognisance   width 1pt $\textbf{69.12}|\textbf{49.31}$ $\textbf{47.64}|\textbf{63.08}$ $\textbf{26.97}|67.01$ $\textbf{53.13}|\textbf{58.16}$   width 1pt $\textbf{60.85}|\textbf{42.30}$ $\textbf{38.46}|\textbf{55.21}$ $\textbf{19.90}|\textbf{60.11}$ $\textbf{44.63}|\textbf{50.78}$
Ensemble TADE   width 1pt $57.30|\textbf{55.22}$ $46.85|50.29$ $\textbf{34.69}|37.93$ $49.21|50.41$   width 1pt $49.61|\textbf{48.19}$ $37.55|42.59$ $\textbf{27.52}|32.21$ $40.87|43.28$
RIDE   width 1pt $63.18|51.44$ $47.67|52.55$ $29.91|47.38$ $51.21|51.33$   width 1pt $54.83|44.02$ $38.25|44.20$ $22.77|38.12$ $42.56|43.21$
@@
* RIDE + Cognisance   width 1pt $\textbf{64.83}|54.60$ $\textbf{50.95}|\textbf{56.21}$ $33.28|\textbf{50.15}$ $\textbf{53.85}|\textbf{54.66}$   width 1pt $\textbf{57.10}|47.63$ $\textbf{42.00}|\textbf{48.42}$ $25.50|\textbf{41.23}$ $\textbf{45.57}|{47.03}$
TABLE II: Evaluation of CLT and GLT protocols on MSCOCO-GLT. The family of ?+Cognisance won for 44 times out of 48 comparisons against the competing models.
@@
Re-balance Baseline   width 1pt $81.27|71.08$ $74.13|76.06$ $50.17|85.61$ $71.88|76.15$   width 1pt $74.59|64.82$ $66.25|69.08$ $35.75|77.22$ $63.10|69.14$
BBN   width 1pt $83.59|70.21$ $76.13|76.30$ $47.25|\textbf{90.98}$ $72.98|77.02$   width 1pt $75.36|61.97$ $68.29|68.61$ $31.75|\textbf{81.38}$ $63.41|68.73$
BLSoftmax   width 1pt $80.77|71.87$ $75.67|69.94$ $45.25|90.08$ $71.31|74.84$   width 1pt $73.23|65.13$ $68.63|62.64$ $32.08|79.60$ $62.81|67.09$
Logit-Adj   width 1pt $81.55|73.95$ $76.00|75.55$ $\textbf{60.83}|82.04$ $74.97|76.28$   width 1pt $73.64|66.98$ $68.71|67.61$ $\textbf{46.58}|70.69$ $66.00|68.01$
GLTv1   width 1pt $82.45|73.09$ $76.42|79.53$ $55.58|86.87$ $74.40|78.60$   width 1pt $76.41|67.07$ $67.33|70.54$ $39.75|81.02$ $65.07|71.39$
@@
* BLSoftmax + Cognisance   width 1pt $83.00|74.17$ $78.37|76.24$ $53.92|86.51$ $75.07|77.58$   width 1pt $75.86|67.98$ $70.46|68.14$ $40.08|80.26$ $66.22|70.59$
* Logit-Adj + Cognisance   width 1pt $82.68|\textbf{77.16}$ $\textbf{80.17}|75.45$ $59.17|83.54$ $\textbf{76.78}|77.77$   width 1pt $75.14|\textbf{70.56}$ $\textbf{71.54}|66.34$ $45.83|73.90$ $\textbf{67.59}|69.50$
Augment Mixup   width 1pt $82.41|72.95$ $73.12|79.53$ $54.33|87.47$ $72.76|78.68$   width 1pt $74.64|66.43$ $65.33|71.00$ $36.00|78.11$ $62.79|70.74$
@@
* RandAug + Cognisance   width 1pt $\textbf{86.55}|\textbf{78.31}$ $81.50|82.82$ $62.42|\textbf{90.29}$ $\textbf{79.47}|\textbf{82.65}$   width 1pt $\textbf{79.05}|69.93$ $\textbf{72.92}|74.51$ $44.33|\textbf{81.82}$ $69.33|74.28$
Ensemble TADE   width 1pt $83.55|76.47$ $80.79|73.89$ $50.50|89.34$ $75.57|78.06$   width 1pt $77.18|68.62$ $71.58|65.63$ $33.50|82.32$ $65.83|70.22$
RIDE   width 1pt $84.23|77.56$ $82.29|77.79$ $58.42|89.69$ $78.09|80.16$   width 1pt $77.45|71.20$ $75.17|68.67$ $41.25|83.33$ $69.02|72.66$
@@
* RIDE + Cognisance   width 1pt $\textbf{86.27}|\textbf{79.64}$ $\textbf{83.13}|\textbf{79.89}$ $\textbf{63.50}|\textbf{90.41}$ $\textbf{80.26}|\textbf{81.98}$   width 1pt $\textbf{80.09}|\textbf{72.77}$ $\textbf{75.17}|\textbf{71.45}$ $\textbf{47.83}|83.03$ $\textbf{71.38}|\textbf{74.34}$
@@ IV-C Comparisons with LT Line-up
TABLE III: Evaluation of ImageNet-GLT and MSCOCO-GLT on LT test set. The family of ?+Cognisance won for 16 times out of 18 comparisons against the competing models.
Benchmarks   width 1pt ImageNet-GLT   width 1pt MSCOCO-GLT
Overall Evaluation   width 1pt $Acc$ $Prec$ $F1$   width 1pt $Acc$ $Prec$ $F1$
BBN   width 1pt 58.60 48.90 53.32   width 1pt 84.84 78.04 81.30
BLSoftmax   width 1pt 51.73 41.97 46.34   width 1pt 83.69 73.81 78.44
Logit-Adj   width 1pt 50.94 40.58 45.17   width 1pt 85.28 73.60 79.01
GLTv1   width 1pt 58.28 50.43 54.07   width 1pt 86.67 81.88 84.21
[…]
Augment Mixup   width 1pt 55.25 46.76 50.66   width 1pt 86.95 82.00 84.40
RandAug   width 1pt 59.88 51.28 55.25   width 1pt 87.64 81.02 84.20
[…]
Ensemble TADE   width 1pt 55.21 46.52 50.49   width 1pt 86.57 79.91 83.10
RIDE   width 1pt 60.17 48.28 53.58   width 1pt 88.03 81.79 84.80
[…]
* RIDE + Cognisance   width 1pt 62.11 50.80 55.89   width 1pt 89.02 81.85 85.28
@@
IV-D Parameter Sensitivity Analysis
Two parameters are used in the construction of CLF: $d_{rn}$ and $d_{rd}$. The parameter $d_{rn}$ represents the radius of the coarse-grained node, while $d_{rd}$ is the radius for density calculation and it can influence the fineness of tree splitting. To facilitate parameter tuning in the experimental session, we introduced the concept of the base distance. The base distance is calculated as the average distance of all sample points to their three nearest neighbors. Then, we use the base distance as the reference value for heuristic parameter tuning, which means setting $d_{rn}$ and $d_{rd}$ as multiples of the base distance instead of blindly adjusting them. As shown in Fig. 10, we fix one parameter and adjust the other parameter. The experimental results indicate that when the parameters are set within a reasonable range, the model performance only exhibits small fluctuations. This indicates that the model is robust to changes in these parameters and can maintain stable performance over a wide range.
(a) Adjustment to $d_{rn}$
(b) Adjustment to $d_{rd}$
Figure 10: Sensitivity analysis of parameters $d_{rn}$ and $d_{rd}$ on ImageNet-GLT, the units of horizontal and vertical coordinates are multiples of the base distance.
IV-E Comparisons with LT Line-up on Noisy Datasets
To test the effectiveness of the noise identification algorithm proposed in this paper, we constructed two long-tailed datasets with noise. These datasets are sampled from Animal-10N and Food-101N, both following a power-law distribution, and the imbalance ratios are 8 and 5, respectively. Detailed parameters of these datasets are shown in Table IV.
TABLE IV: Detailed information of Animal-10NLT and Food-101NLT datasets
Dataset Name #Training #Test #Classes Noise Rate (%)
Animal-10NLT 19,261 5,000 10 $\simeq$8
Food-101NLT 19,230 3,824 101 $\simeq$20
Animal-10N 50,000 5,000 10 $\simeq$8
Food-101N 52,868 3,824 101 $\simeq$20
TABLE V: Performance comparison of various methods on Animal-10NLT and Food-101NLT datasets
Dataset   width 0.75pt Animal-10NLT   width 0.75pt Food-101NLT
@@ IV-E Comparisons with LT Line-up on Noisy Datasets
Long-Tail Classification Baseline   width 0.75pt 58.70 61.64 60.14   width 0.75pt 21.30 22.49 21.88
BLSoftmax   width 0.75pt 59.72 60.59 60.15   width 0.75pt 20.02 22.17 21.04
Logit-Adj   width 0.75pt 60.04 60.76 60.40   width 0.75pt 21.94 21.56 21.75
Mixup   width 0.75pt 61.94 67.98 64.82   width 0.75pt 27.02 30.86 28.81
RandAug   width 0.75pt 67.18 69.40 68.27   width 0.75pt 29.09 30.71 29.87
TADE   width 0.75pt 63.80 64.27 64.03   width 0.75pt 26.28 30.28 28.14
RIDE   width 0.75pt 70.52 70.01 70.26   width 0.75pt 28.62 30.12 29.35
@@
Co-Teaching+   width 0.75pt 56.19 54.23 55.19   width 0.75pt 9.96 6.01 7.5
RandAug + Co-Teaching   width 0.75pt 57.03 62.33 59.56   width 0.75pt 17.8 18.12 17.96
RandAug + Co-Teaching+   width 0.75pt 51.93 41.68 46.24   width 0.75pt 8.54 3.82 5.28
@@
* Baseline + Cognisance+   width 0.75pt 69.05 72.50 70.72   width 0.75pt 31.78 37.12 34.22
* BLSoftmax + Cognisance+   width 0.75pt 70.12 72.83 71.45   width 0.75pt 30.94 35.84 33.25
* Logit-Adj + Cognisance+   width 0.75pt 71.30 73.55 72.41   width 0.75pt 31.25 36.11 33.50
* TADE + Cognisance+   width 0.75pt 73.82 74.70 74.26   width 0.75pt 33.02 37.78 35.23
* RIDE + Cognisance+   width 0.75pt 73.20 74.12 73.65   width 0.75pt 33.46 38.35 35.73
[…]
Compared to the Cognisance framework proposed in the previous chapter, the Cognisance+ framework shows significant performance improvement on the Food-101NLT dataset, which has high noise rates and fewer samples per class. Additionally, we observed that the Co-Teaching series performs worse than the Baseline on long-tail noisy datasets. This is because its noise filtering method relies on confidence, and samples from tail classes, due to their rarity, usually have low confidence, leading to a large number of misclassifications.
Since our framework already incorporates the RandAug method, we choose not to integrate it with other data augmentation methods like Mixup. Moreover, Co-Teaching methods require maintaining two models simultaneously, and our experimental setup also necessitates maintaining two models. Although our framework is capable of integrating with the Co-Teaching series, we decided against it to avoid overly complex models. Furthermore, we found convergence issues with the Co-Teaching+ method on the Food-101NLT datasets. Despite multiple random seed adjustments, its performance remained poor. Since Co-Teaching+ is based on the Co-Teaching method, and considering that Co-Teaching’s performance on these datasets, though low, is within a normal range, we speculate that this issue may be related to the “disagreement” samples training used by Co-Teaching+. When the performance of the two original Co-Teaching models is low, the two teacher models may have reached a consensus in their errors.
@@ V-A Why not directly apply the OLeaF?
In Cognisance, CLF is more appropriate compared to OLeaF for two very important characteristics: 1) Automatic tree-splitting mechanism. The number of parameters of CLF is relatively small, and the fine degree of tree-splitting can be controlled only by adjusting the parameter $d_{rd}$. The tree-splitting scheme in OLeaF is more rigorous and detailed but requires manually tuning parameters $d_{c}$ and number of trees. However, the feature learner needs to iterate in this framework and it may be possible to re-trigger clustering after each iteration. Besides, the target datasets for Cognisance are often large image data, so a fully automated scheme is compulsory; 2) Computing efficiency and representativeness. Each node of CLF is a coarse-grained node because the head attribute may contain a large number of similar samples. If left un-merged, the length of the path occupied by the head attribute may skyrocket. While each node in the same path is sampled with the same weight, a large number of extremely similar nodes actually lose their representativeness. At the same time, the sampling weights of bottom nodes are compromised. So it is necessary to carry out a small range of clustering to be controlled, and this clustering process is mainly controlled by the parameter $d_{rn}$. All the member nodes within a coarse node will equally share the node’s sampling weight.